LaurenFisher:

The equation d = d equals StartFraction m Over V EndFraction. can be used to calculate the density, d, of an object with mass, m, and volume, V. Which is an equivalent equation solved for V?

supie:

Is there answer choices?

PWRDrEaMz:

Look at the choices carefully and maybe read the question over

LaurenFisher:

dm = V d/m = V m/d= V m – d = V

snowflake0531:

Oh wait, sorry that would be a direct answer. I only meant, you could search the relationship between mass, volume, and density

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so given \[d = \frac{ m }{ V }\] using this equation solve it for V ,so in this way will get the right choice ... \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

@supie